3. Determine the value of the equilibrium constant, Q for the chemical reaction given below, if 0.108 moles of H2, 0.087 of Iodine, 0.020 of Ni , Si with 0.110 and water vapor are in a 6 liters drum and at an equilibrium.


H2= 0.108


I= 0.087


Ni= 0.020


Si= 0.110


6 liters


substitute the value of the compounds


1
Expert's answer
2021-04-05T07:56:54-0400

volume of solution = 6L

moles of H2 = 0.108 moles

Concentration of H2 = 0.108/6 = 0.018

moles of iodine I2 = 0.087 moles

Concentration of I2 = 0.087/6 = 0.0145

moles of Ni = 0.020 moles

Concentration of Ni = 0.020/6 = 0.0033

moles of Si = 0.010 moles

Concentration of Si = 0.010/6 = 0.0016


Now equilibrium constant = "\frac{[H_2][I_2]}{[Ni][Si]}"

= "\frac{[0.018][0.145]}{[0.0033][0.0016]}" = 49.43